find the slope, if it exists, of the line containing the pair of points.
(7,8) and (9,-7)
a.the slope m=?
b. the slope is undefined

Answer by stanbon(75887)

find the slope, if it exists, of the line containing the pair of points.
(7,8) and (9,-7)
a.the slope m= (-7-8)/(9-7) = -15/2 =
-------
b. the slope is undefined: No
=================================
Cheers,
Stan H.
=========